OpenStudy (anonymous):

The half-life of carbon-14 is 5,730 years. How many grams of parent material will remain from a 120-gram sample of carbon-14 after 2 half-lives? 90 60 30 15

OpenStudy (unklerhaukus):

after each half life the amount of C-12 decreases by half after two half-lives the amount of C-12 has decreased by half twice,

OpenStudy (anonymous):

Thanks

OpenStudy (unklerhaukus):

\[120\times\frac 12\times \frac 12\]

OpenStudy (anonymous):

I am looking at the answer as 30

OpenStudy (unklerhaukus):

good good

OpenStudy (unklerhaukus):

a 120g sample of pure C-14 will be 30g after 11460 years

OpenStudy (unklerhaukus):

*i mean C-14 in my first post not C-12. C-12 is radioactively stable.